Tuesday, July 6, 2010

Summary Night Report 20100706

Report for night 2010-07-06

SA: Alexei
SO: Fred
Others: Hitesh, Deon, Siphelo, Zolisa

Notes and messages


1. SPS error was generally fixed by Hitesh and Deon during day time.
Segments 3 and 41 were disconnected.

2. Two new SOs arrived today and Fred started teaching.

3. The alignment is working in general, but we need good conditions
to make it properly.

No comments:

Post a Comment